Consider a wind turbine having moment of inertia 240,000 kg m2. If a 120 kg person were holding on to the edge of the turbine 10 m from its center the torque due to the person is ________ and the angular acceleration of the turbine is ________.

120,000 Nm; 0.05 rad/s2

12,000 Nm; 0.005 rad/s2

240,000 Nm; 0.05 rad/s2

12,000 Nm; 20 rad/s2
Answer #1

Torque formula

τ = F * d = 120kg * 10m/s² * 10m = 12 000 N·m (so eliminate 1 and 3)
also τ = 12000 N·m = Iα = 240000kg·m² * α

after solving
α = 0.05 rad/s²